OpenStudy (anonymous):

The diagonals of rhombus FGHJ intersect at point K. If side GH is equal to x + 9 and side JH is equal to 5x – 2, find x.

OpenStudy (anonymous):

so 5x-2 = x+9 4x = 11 x = 11/4 = 2.75

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Thankss!!